OpenStudy (anonymous):

Prove, algebraically and logically, that the combination equation: nCr + nC(r+1) = (n+1)C(r+1)

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ r\end{matrix}\right) + \left(\begin{matrix}n \\ r+1\end{matrix}\right) = \left(\begin{matrix}n+1 \\ r+1\end{matrix}\right)\]

OpenStudy (anonymous):

Do you mean algebraic and combinatorial?

OpenStudy (anonymous):

Just without the use of number substitutions.

OpenStudy (anonymous):

Why don't you just expand it and see ...

OpenStudy (anonymous):

How are you meant to expand it. My teacher showed me very quickly today then rubbed it out, all I remember is it started by changing it to factorial form.

OpenStudy (anonymous):

\( \huge \binom{n}{r} = \frac{n!}{r!\times(n-r)!} \)